OpenStudy (anonymous):

Please Help!! How many 3-digit codes using the digits 0-9 are possible if repetitions are allowed?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

But how did you know what to multiply?

OpenStudy (anonymous):

Ok, just getting into that now. Thank you

OpenStudy (anonymous):

Luis Rivera's answer seems incorrect (as repetition is allowed) the correct answer is \( 9 \times 10 \times 10 = 900 \).

OpenStudy (zarkon):

are leading zeros permitted?

OpenStudy (anonymous):

I assumed leading zeros are not permitted since in general sense that will not be a 3 digit number.

OpenStudy (zarkon):

but it says 3 digit code not 3 digit number

OpenStudy (zarkon):

if you allow leading zeros then there are \(10^3=1000\) possibilities

OpenStudy (anonymous):

Okay, if the leading zero is allowed then you have to add 100 more numbers to the previous 900.

OpenStudy (anonymous):

\[ 10 \times 10 \times 10 \]

OpenStudy (zarkon):

@Luis Rivera ?

OpenStudy (anonymous):

Yes, that last comment doesn't make any sense.